### v1

# On the Limit of Language Models as Planning Formalizers

Abstract
Large Language Models have been shown to fail to create executable and verifiable plans in grounded environments. An emerging line of work shows success in using LLM as a formalizer to generate a formal representation (e.g., PDDL) of the planning domain, which can be deterministically solved to find a plan. We systematically evaluate this methodology while bridging some major gaps. While previous work only generates a partial PDDL representation given templated and thus unrealistic environment descriptions, we generate the complete representation given descriptions of various naturalness levels. Among an array of observations critical to improve LLMs’ formal planning ability, we note that large enough models can effectively formalize descriptions as PDDL, outperforming those directly generating plans, while being robust to lexical perturbation. As the descriptions become more natural-sounding, we observe a decrease in performance and provide detailed error analysis. [^1] Our code and data can be found at https://github.com/CassieHuang22/llm-as-pddl-formalizer.
On the Limit of Language Models as Planning Formalizers
Cassie Huang  Li Zhang Drexel University {Cassie.Huang}@drexel.edu{Harry.Zhang}@drexel.edu
1 Introduction
Large language models (LLMs) can make informal plans, such as suggesting ideas for parties or giving general advice on immigration. However, most users, let alone automated agents like robots, would not be able to actually execute those plans step-by-step to fruition – either to organize parties or acquire visas – without significant prior knowledge or external help. This inability to make executable plans lies in LLMs’ inability of grounding and formal reasoning Liu et al. (2023b); Pan et al. (2023); Zhang et al. (2023). Cutting-edge research in the community has evaluated LLMs’ ability to make formal plans in grounded environments, such as textual simulations, where all objects and actions represent actualities in the real world. Therefore, any resulting plan that formally involves those objects and actions would be executable and verifiable by nature. Although formal planning has been desirable in the history of AI Weld (1999), recent work found that even state-of-the-art LLMs are unable to generate formal plans Silver et al. (2024); Valmeekam et al. (2024); Stechly et al. (2024).
Figure 1: LLM-as-formalizer uses natural language descriptions to generate the Domain and Problem File in PDDL, then these are given to a planner to find a plan. We explore the effect of natural-ness of the language in the description, by giving the model both templated and natural descriptions. Examples of Domain Descriptions and Problem Descriptions from the Blocksworld Domain are shown. The green text displays what the two examples have in common (listing all possible actions and restrictions) and the red text displays text that is not considered natural. The “Templated” text corresponds to the “Heavily Templated” version discussed in Section 4.
Instead of using the LLM as a planner to generate the plan directly, an alternative line of work uses the LLM as a formalizer. Here, the LLM generates a formal representation of a planning domain, for example in the planning domain definition language (PDDL), based on some natural language descriptions of the environment. This representation can then be fed into a solver to find the plan deterministically (see Figure 1). Previous work achieved great success by showing that LLM-as-formalizer greatly outperforms LLM-as-planner in various domains Lyu et al. (2023); Xie et al. (2023); Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024); Zhang et al. (2024c); Zhu et al. (2024), as LLMs are more capable of information extraction than formal reasoning Zhang et al. (2024b). However, the above work has two major shortcomings. First, to simplify the task and evaluation, most have only attempted to generate a partial PDDL representation while assuming the rest is provided, often unrealistic in real life. Second, the language used to describe the environments is often artificially templated and structured, leading to potential overestimation of models’ ability.
This paper explores the limit of LLM-as-formalizer devoid of the above two simplifications. We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1). On one of the most widely used planning datasets, BlocksWorld PIC (1998) and its derivative, MysteryBlocksWorld Valmeekam et al. (2024), we benchmark a suite of LLMs on generating PDDL that is both solvable and correct. As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
Our work is the first to systematically analyze state-of-the-art LLMs’ ability of the trending methodology of LLM-as-formalizer on the highly challenging task of formal planning. We put forward an array of observations that will benefit future efforts. Discussed in detail in Section 5, our key findings are as follows.
  • On fully-observed environments such as the BlocksWorld domain, larger GPT models can decently generate entire PDDL, while smaller open-source models cannot.
  • When feasible, LLM-as-formalizer greatly outperforms LLM-as-planner.
  • As the environment descriptions sound more human-like, the models are more challenged.
  • The performance of LLM-as-formalizer is robust to lexical perturbation, while that of LLM-as-planner is not.
  • Errors in PDDL generation span syntax and semantics in both the domain and problem.
2 Task: Formal Planning with PDDL
Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1). $\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance. Concretely, the $\mathbb{DF}$ defines all available actions for a state-based environment as well as predicates that represent the properties of object types. Each action definition contains the name of the action, parameters and semantics. The semantics of an action include the preconditions which describe the necessary world states where the action is valid to execute, and effects which describe how the states change after the action is executed. The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states. These two files are then given to a deterministic planner which will algorithmically search for a plan. Such a plan is a series of executable, instantiated actions that sequentially transition the world states from initial to goal. In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
However, formal planning demands a well-crafted pair of $\mathbb{DF}$ and $\mathbb{PF}$. In a real-world planning scenario, an average user would not describe their environment and problem with PDDL, but more likely with a textual description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{PD}$), which can be specific or loose. Hence, we focus on the textual flavor of formal planning: given $\mathbb{DD}$ and $\mathbb{PD}$, the model outputs a successful plan with regard to the $\mathbb{DF}$ and $\mathbb{PF}$ that are withheld from the model.
3 Methodology: LLM-as-Formalizer
Figure 2: Methodologies for using LLMs in Formal Planning. LLM-as-Planner generates the plan directly, while LLM-as-Formalizer translates the $\mathbb{DD}$ and $\mathbb{PD}$ into PDDL. Previous work such as Liu et al. (2023a) use the LLM to generate partial PDDL, while we generate the entire PDDL (dashed arrow).
To tackle the task above, recent work involving LLMs diverged into two methodologies. The first, LLM-as-planner, directly uses LLMs to generate a plan based on the $\mathbb{DD}$ and a $\mathbb{PD}$. The second, LLM-as-formalizer, uses LLMs to recover the $\mathbb{DF}$ and $\mathbb{PF}$, before using a deterministic planner to arrive at the plan (Figure 2). Our work will focus on the second while using the first as a baseline. LLM-as-formalizer is in essence neurosymbolic, where LLMs help define the structured representation that is otherwise prohibitively costly to annotate and brittle to generalize. Existing works in this line demonstrated success but only generated a partial PDDL representation, while assuming the rest, including $\mathbb{PF}$ goals Lyu et al. (2023); Xie et al. (2023), the $\mathbb{PF}$ Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024), the action semantics in the $\mathbb{DF}$ Zhang et al. (2024c); Zhu et al. (2024), and the domain file Wong et al. (2023); Guan et al. (2023). While this simplifies the task and evaluation, the assumption of provided PDDL components is often unrealistic. Therefore, we bridge this gap by asking the LLM to predict the entire PDDL – both the $\mathbb{DF}$ and $\mathbb{PF}$ [^2] It is however minimally necessary to provide the action space, the identifiers and parameters of the actions in $\mathbb{DF}$, so the agent knows what actions are possible.
4 Evaluation: Datasets, Models, Metrics
To evaluate both approaches above, we work with fully-observed textual environments. Here, the provided $\mathbb{DD}$ and $\mathbb{PF}$ contain all necessary information for the model to make a complete plan.
4.1 Data
We consider two datasets, BlocksWorld PIC (1998) and MysteryBlocksWorld Valmeekam et al. (2024), one of the most widely used domains in related work. Here, the model must rearrange stacks of blocks on a table from an initial configuration to a goal configuration using a single arm. While each instance has ground-truth $\mathbb{DF}$ and $\mathbb{PF}$, they are invisible to the models and for evaluation only. Notwithstanding existing datasets, we generate the problem configurations ourselves to maximize control over ablation studies (see Section 5). Problem configurations were generated by randomly sampling the number of blocks between 2 and 15 and number of stacks in the initial state and goal between 1 and the number of blocks. Then, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created using 3 methods for 3 different levels of ‘naturalness’.
Heavily Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al. (2024). This description is almost a word-by-word translation of PDDL. For example, for the ‘pick-up’ action, the ground-truth PDDL $\mathbb{DF}$ would be the following:
(:action pick-up :parameters (?b - block) :precondition (and (clear ?b) (on-table ?b) (arm-empty)) :effect (and (not (on-table ?b)) (not (clear ?b)) (not (arm-empty)) (holding ?b)) )
while the Heavily Templated $\mathbb{DD}$ is:
To perform Pickup action, the following facts need to be true: clear block, block on table, arm-empty. Once Pickup action is performed the following facts will be true: holding block. Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty.
From an application point of view, spelling out all preconditions and effects in terms of the predicates is paradoxical, as it assumes the user already have the algorithmic awareness of PDDL.
Moderately Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as the original BlocksWorld dataset, following Valmeekam et al. (2024). For example. the Moderately Templated description of the ‘pick-up’ action is:
I can only pick up or unstack one block at a time. I can only pick up or unstack a block if my hand is empty. I can only pick up a block if the block is clear. A block is clear if the block has no other blocks on top of it and if the block is not picked up.
While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’.
Natural. A realistic pair of $\mathbb{DD}$ and $\mathbb{PD}$ should emulate how real-life users would describe the planning domain and problem, such that a human problem solver would understand and have just enough information to find a plan. To create such descriptions, we use a human-in-the-loop, model-assisted data generation approach.
To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$ for BlocksWorld, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique. We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
To generate $\mathbb{PD}$, we first algorithmically and randomly permutate BlocksWorld problems, given various total numbers of blocks and stacks. We then provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration. The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars. We manually verify the correctness of the dataset of the non-templated problems by hand by comparing them against the problem configurations. We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
The robot arm can pick up and move one block at a time from one position to another. It is only able to move the top block from any stack or table, and have only one block held by the robot arm at a time. The main actions available are ’pick up’, …
The above example of the Natural description no longer discusses the preconditions and effects of each actions one by one, but rather focuses on the general rules to the domain. These rules apply to not only ‘pick-up’ but also other actions. Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and more realistic.
In total, we construct 111 problems each varying in the number of objects in the unified domain of BlocksWorld. For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 111 $\mathbb{PD}$s. For the Natural description, there are 111 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s. We refer to this dataset as BlocksWorld-111.
Mystery BlocksWorld is a dataset created in Valmeekam et al. (2024) that obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test (see an example in Appendix A). This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization. The original Mystery BlocksWorld dataset is Heavily Templated. Since the dataset itself is an artificial perturbation, we do not provide more natural descriptions as that would defeat its purpose. We sampled 100 random problems for our experiment, resulting in 1 $\mathbb{DF}$, 1 $\mathbb{DD}$, 100 $\mathbb{PF}$s, and 100 $\mathbb{PD}$s. We refer to this dataset as MysteryBlocksWorld-100.
Data examples can be found in Appendix A.
4.2 Metrics
Following past work Guan et al. (2023); Zhu et al. (2024), the model-predicted plan is validated using VAL Howey et al. (2004) against the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ provided above, instead of being compared against “ground-truth” plans like some work Lyu et al. (2023); Liu et al. (2023b); Pan et al. (2023) since there could be multiple correct plans. For the LLM-as-formalizer approach, the predicted $\mathbb{DF}$ and $\mathbb{PF}$ are similarly not compared against the ground-truth, as only the eventual plan is validated because there might be more than one way to formalize the planning domain and problem in PDDL.
We evaluate the predicted plans following Zuo et al. (2024): solvability and correctness. Solvability indicates the percentage of plans that were able to be found based on the predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether it can correctly work with the actual environment (in our case, described by the gold $\mathbb{DF}$ and $\mathbb{PF}$). In contrast, correctness is a subset that indicates the percentage of actually correct plans. Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
4.3 Models
For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b-it, gemma-2-27b-it Team et al. (2024), llama-3.1-8B-Instruct, llama-3.1-70B-Instruct Dubey et al. (2024), gpt-3.5-turbo-0125, gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o1-preview-2024-09-12444platform.openai.com/docs/models o1 is only reported in some experiments due to prohibitive cost.. We query these models using KANI Zhu et al. (2023) with default hyper-parameters. The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-111. To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
5 Results and Observations
In this section, we display our results as well as perform an in-depth analysis of the strengths and weaknesses of LLMs in formal planning, to understand the impact of the model choice, naturalness of the description, content of the task, and difficulty of the problem.
5.1 Can LLMs formalize?
Models Solvability Correctness
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gemma-2-9b-it 3/111 3/111
gemma-2-9$\dagger$ - 10/111
gemma-2-27b-it 0/111 0/111
gemma-2-27$\dagger$ - 12/111
gpt-3.5-turbo 2/111 1/111
gpt-4o-mini 22/111 3/111
gpt-4$\dagger$ - 7/111
gpt-4o 73/111 66/111
gpt-4$\dagger$ - 37/111
o1-preview 101/111 101/111
o1$\dagger$ - 93/111
Table 1: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Natural BlocksWorld-111. Similarly trends hold for the Heavily and Moderately Templated versions (results are shown in Appendix C).
First, we seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work. Table 1 displays the results on our experiment using the most natural sounding descriptions on the BlockWorld.
These results demonstrate that big enough LLMs can decently generate PDDL on fully observed environments. We can see that for BlocksWorld, the performance for gpt-4o surpasses the performance for gpt-4o-mini, which surpasses the performance for gpt-3.5-turbo. As the model gets larger, the number of syntax errors decreases, indicating larger models’ stronger ability of code generation. This conclusion also holds for the other levels of naturalness (Figure 3).
We also observe that open-sourced models with size up to 70B almost cannot generate PDDL. Llama models cannot able to find a single plan across all three datasets while gemma models show poor though non-zero performance across all three datasets. Detailed analysis of errors will be discussed in Section 5.5.
5.2 Should LLMs formalize?
Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology for symbolic planning tasks like BlocksWorld? From Table 1, we can see that the best performing gpt-4o is able to generate solvable PDDL 73/111 times, and of those 73 plans, 66 of them are correct. This far surpasses the LLM-as-planner baseline, which only found correct plans 37/111 times. This conclusion holds for the Templated BlocksWorld-111 data (Figure 3). From Figure 3, for MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox. For gpt-4o-mini, LLM-as-formalizer still outperforms LLM-as-planner for all cases (all naturalness level across two datasets) except Natural BlocksWorld-111.
These results demonstrate that LLMs as formalizers greatly outperforms those as planners in most cases, whenever these LLMs can foramlize PDDL at all. However, these results also show that models that might not be able to formalize (eg. Llama models) but might still be able to plan, though they will have low performance.
5.3 The more natural, the harder?
Figure 3: Performance across different naturalness level of description on BlocksWorld-111.
In this section, we discuss the question of whether using humanized descriptions makes the problem more difficult. Results from Figure 3 show that as the problem sounds more similar to PDDL, and less natural, the performance of all the models improves. Also, from our results in the Mystery BlocksWorld domain, LLM-as-formalizer performs quite well for a few of the models. However, looking at the domain and problem descriptions, they are very templated and sound like PDDL. This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions. One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage. Another reason is a more-natural sounding description may leave out implicit common-sense. For example, the Natural BlocksWorld-111 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved. However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
5.4 Do LLMs memorize pretraining?
Do LLMs generate plans or formalize PDDL based on what they have memorized in their training data? We determine this by looking at the results on MysteryBlocksWorld-100, a derivative of BlocksWorld where all names are perturbed and nonsensical. From Figure 3, we can see that LLM-as-planner was not able to find a single correct plan using either gpt-4o-mini or gpt-4o. However, gpt-4o-as-formalizer surpassed this baseline with a Correctness score of 70/100. This suggests that LLM-as-formalizer is robust to lexical perturbation, and its success is not due to memorization of the domain which is a part of the pretraining data.
5.5 What kind of errors?
Natural BlocksWorld-111
Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
gemma-2-9b-it 15/20 20/20 20/20
gemma-2-27b-it 3/20 20/20 14/20
Llama-3.1-8B 20/20 20/20 18/20
Llama-3.1-70B 20/20 20/20 17/20
gpt-3.5-turbo 10/20 20/20 20/20
gpt-4o-mini 2/20 20/20 19/20
gpt-4o 2/20 2/20 18/20
Table 2: Different model errors in generating PDDL for all the natural BlocksWorld-111, manually annotated on a 20–example subset.
MysteryBlocksWorld-100
Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
gpt-3.5-turbo 16/20 19/20 2/20
gpt-4o-mini 6/20 20/20 1/20
gpt-4o 5/20 16/20 0/20
Table 3: Different errors of gpt-3.5-turbo, gpt-4o-mini and gpt-4o on Heavily Templated MysteryBlocksWorld-100, manually annotated on a 20–example subset.
Natural BlocksWorld-111
Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
gpt-4o-mini 11/20 18/20 19/20 1/20 2/20
gpt-4o 0/20 2/20 0/20 0/20 0/20
Heavily Templated MysteryBlocksWorld-100
Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
gpt-4o-mini 14/20 17/20 17/20 0/20 5/20
gpt-4o 13/20 14/20 0/20 0/20 2/20
Table 4: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-111 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.
In this section, we discuss the kind of errors in PDDL generation. We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct. From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$. Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters. The error analysis can be found for the Natural BlocksWorld-111 in Table 2 and Table 4.
As mentioned before, as the models get larger, the amount of syntax errors in the generated PDDL decreases. Smaller models consistently make mistakes in generating the $\mathbb{DF}$. Interestingly, the most common error made for gpt-4o came from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$. Common errors in the $\mathbb{PF}$ included incorrect predicates in the initial state and goal state, possibly because the descriptions are natural rather than templated, so that the implicit information in the $\mathbb{PD}$ may be nontrivial to infer. Regarding the $\mathbb{DF}$, the most common error is an incorrect effect in an action. For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand.
For the open-source models, the most common error is syntax errors. For examples, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL language. The smallest gemma model also made many mistakes in syntax, though fewer than the Llama models, However, despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates and incorrect effects in the actions. For the Heavily Templated BlocksWorld-111 results, there is a significant gap between the number of plans that were found, and the number of found that were correct. Upon looking at the generated PDDL, we found that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans.
For the MysteryBlocksWorld-100 dataset (Table 3), we see a similar trend of the syntax errors decreasing as the models get larger. For all three GPT models analyzed, there are barely any errors in the $\mathbb{PF}$ but rather the most common errors came from the $\mathbb{DF}$. Since the Mystery BlocksWorld domain is a result of lexical perturbation, the task of PDDL generation is akin to symbolic information extraction or translation, devoid of much use of commonsense knowledge. As the descriptions for MysteryBlocksWorld-100 were heavily templated, sounding the most similar to PDDL, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$. While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved. From Table 4, a similar trend between BlocksWorld-111 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.
6 Related Work
Planning with LLMs There has been a large amount of research using LLMs for planning tasks. Some use LLMs for informal planning, also known as script or procedure learning Zhang et al. (2020); Lyu et al. (2021); Lal et al. (2024). While modern LLMs can make coherent and plausible informal plans, they are ungrounded and so lack executability and verifiability. Work that use LLMs for formal planning in grounded environments generally conclude the inability of such LLMs-as-planners Silver et al. (2024); Valmeekam et al. (2024); Stechly et al. (2024). Follow-up work tackles this shortcoming by using the LLM as a heuristic, not just a planner, such as by proposing candidate plans that are iteratively verified Valmeekam et al. (2023); Kambhampati et al. (2024). While we consider the standard LLM-as-planner as a baseline, our focus is on LLM-as-formalizer, an alternative methodology for the same problem.
LLMs as PDDL Formalizer Here, LLMs do not provide plans but rather generate the a PDDL representation of the domain and problem, which is then run through a solver to find the plan. This methodology has proven successful in a number of recent works, where the LLM generates different parts but not all of the PDDL for simplified evaluation. Zuo et al. (2024); Zhang et al. (2024a); Liu et al. (2023a) use the LLM to predict the entire $\mathbb{PF}$, while Xie et al. (2023); Lyu et al. (2023) predict just the goal for the $\mathbb{PF}$. Some predict the $\mathbb{DF}$, such as Zhang et al. (2024c); Zhu et al. (2024) that generate the action semantics of the $\mathbb{DF}$ and Wong et al. (2023) who also predicts the predicates from a candidate list. Closest to our work is Guan et al. (2023) which predicts the $\mathbb{DF}$ as well as the $\mathbb{PF}$ goal. However, our work of holistically generating PDDL shows that coming up with the initial state in the $\mathbb{PF}$ is non-trivial (Section 5.5). Moreover, we vary the level of naturalness of descriptions in addition to the templated ones, which prove to be more challenging and insightful (Section 5.3).
While the above discussions pertain to LLMs generating PDDL, many work on embodied agents outside the NLP community tackle similar problems with different focus Li et al. (2024).
LLMs as Formalizer Our work hinges on modern LLMs’ ability to generate code Chen et al. (2021). In addition to writing or debugging programs Jiang et al. (2024), LLMs are also used to generate formal, interim representations that are not necessarily PDDL for problem solving. For example, Gao et al. (2023); Lyu et al. (2023); Tang et al. (2024) use the LLM to generate executable Python code for solving symbolic problems. In other work, the generated code may not be executable and is provided to another LLMs to facilitate reasoning Madaan et al. (2022); Zhang et al. (2023).
A table comparing a couple of these works can be seen in Table 7 in the Appendix (Section F).
7 Conclusion
We explore the limit of state-of-the-art LLMs to be used as a formalizer by generating the entire $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL given natural language descriptions of different naturalness levels. We find that the performance depends on a number of factors, including the size of the model and the naturalness of the descriptions. While the LLM-as-formalizer methodology greatly outperforms the LLM-as-planner methodology in domains exemplified by BlocksWorld, we conclude that with zero-shot prompting only large GPT models are currently capable for the task. Therefore, future work should attempt to equip open-source models with similar ability to democratize the ability of making executable plans. We also find that LLM-as-formalizer is robust to lexical perturbation, demonstrating strong performance in long-tail domains that are not represented in LLMs’ pretraining. Our work will hopefully drive future work on improving LLM as a planning formalizer,including experiments on partially-observed environments that require exploration and interaction, more complex environments with a larger action space, and so on.
8 Limitation
While BlocksWorld is one of the most frequently used domains in LLM’s formal planning, it is the only domain considered in this work, due to the choice to optimize for the analysis’ depth instead of breadth. As future work, we are currently considering complex domains such as Barman in IPC. Even so, a common and valid criticism for using those simulations or text problems for evaluation is that these settings may be too contrived and removed from the reality. Nevertheless, it is likely that LLMs’ satisfactory performance on these datasets is a necessary condition to success in real life.
While we only consider zero-shot prompting without any attempt for prompt tuning, it is possible that the models’ performance significantly increases otherwise. Therefore, experimental results in all settings may be underestimated. Moreover, advanced prompting techniques such as chain-of-thought, self-refine, and voting can all potentially improve model performance. However, the study of those is out of the scope of this work.
While we advocate for the LLM-as-formalizer methodology over LLM-as-planner, the former’s success may be dependent on the task. Highly symbolic tasks which can be relatively easily described, like BlocksWorld, are likely to favor LLM-as-formalizer. However, LLM-as-planner might shine in tasks with a more complex action space requiring common-sense knowledge that is easily accessed by pretraining. Furthermore, while we only consider the most straightforward LLM-as-planner prompting method, more involved methods, like Kambhampati et al. (2024) that combines LLM-as-planner with symbolic validation, will likely lead to a stronger baseline.
Since this work uses only the BlocksWorld and Mystery BlocksWorld domains, it is a small toy example to the usage of LLMs as formalizers and are not representative to problems in the real world, which would be much more challenging. This may pose a risk to users using this code on real world problems.
The datasets we use and we propose are all under the MIT License.
Appendix A Data Examples
A.1 BlocksWorld-111 PDDL
The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for BlocksWorld-111.
$\mathbb{DF}$:
(define (domain blocksworld) (:predicates (clear ?x) (on-table ?x) (arm-empty) (holding ?x) (on ?x ?y)) (:action pickup :parameters (?ob) :precondition (and (clear ?ob) (on-table ?ob) (arm-empty)) :effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob)) (not (arm-empty)))) (:action putdown :parameters (?ob) :precondition (holding ?ob) :effect (and (clear ?ob) (arm-empty) (on-table ?ob) (not (holding ?ob)))) (:action stack :parameters (?ob ?underob) :precondition (and (clear ?underob) (holding ?ob)) :effect (and (arm-empty) (clear ?ob) (on ?ob ?underob) (not (clear ?underob)) (not (holding ?ob)))) (:action unstack :parameters (?ob ?underob) :precondition (and (on ?ob ?underob) (clear ?ob) (arm-empty)) :effect (and (holding ?ob) (clear ?underob) (not (on ?ob ?underob)) (not (clear ?ob)) (not (arm-empty)))))
The $\mathbb{DF}$ contains all four actions (pickup, putdown, stack and unstack) and their pre-conditions and post-conditions, as well as predicates needed for the domain.
$\mathbb{PF}$:
(define (problem blocksworld-p99) (:domain blocksworld) (:objects red blue green yellow ) (:init (on-table red) (on blue red) (clear blue) (on-table green) (on yellow green) (clear yellow) (arm-empty) ) (:goal (and (on-table red) (on green red) (on yellow green) (on blue yellow) )) )
The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
A.2 BlocksWorld-111 $\mathbb{DD}$ and $\mathbb{PD}$
The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the BlocksWorld-111 dataset. We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL. The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
Heavily Templated $\mathbb{DD}$:
I am playing with a set of blocks. Here are the actions I can do Pickup block Unstack block from another block Putdown block Stack block on another block I have the following restrictions on my actions: To perform Pickup action, the following facts need to be true: clear block, block on table, arm-empty. Once Pickup action is performed the following facts will be true: holding block. Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty. To perform Putdown action, the following facts need to be true: holding block. Once Putdown action is performed the following facts will be true: clear block, block on table, arm-empty. Once Putdown action is performed the following facts will be false: holding block. To perform Stack action, the following needs to be true: clear block2, holding block1. Once Stack action is performed the following will be true: arm-empty, clear block1, block1 on block2. Once Stack action is performed the following will be false: clear block2, holding block1. To perform Unstack action, the following needs to be true: block1 on block2, clear block1, arm-empty. Once Unstack action is performed the following will be true: holding block1, clear block2. Once Unstack action is performed the following will be false:, block1 on block2, clear block1, arm-empty.
Heavily Templated $\mathbb{PD}$:
As initial conditions I have that, the blue block is clear, the yellow block is clear, arm-empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table. My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
For the Moderately Templated data, the $\mathbb{DD}$ and $\mathbb{PD}$ are much more natural than the Heavily Templated data, but all predicates are still listed.
Moderately Templated $\mathbb{DD}$:
I am playing with a set of blocks where I need to arrange the blocks into stacks. Here are the actions I can do Pick up a block Unstack a block from on top of another block Put down a block Stack a block on top of another block I have the following restrictions on my actions: I can only pick up or unstack one block at a time. I can only pick up or unstack a block if my hand is empty. I can only pick up a block if the block is on the table and the block is clear. A block is clear if the block has no other blocks on top of it and if the block is not picked up. I can only unstack a block from on top of another block if the block I am unstacking was really on top of the other block. I can only unstack a block from on top of another block if the block I am unstacking is clear. Once I pick up or unstack a block, I am holding the block. I can only put down a block that I am holding. I can only stack a block on top of another block if I am holding the block being stacked. I can only stack a block on top of another block if the block onto which I am stacking the block is clear. Once I put down or stack a block, my hand becomes empty. Once you stack a block on top of a second block, the second block is no longer clear.
Moderately Templated $\mathbb{PD}$:
As initial conditions I have that, the blue block is clear, the yellow block is clear, the hand is empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table. My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
Finally for the natural data, we can see that the $\mathbb{DD}$ and $\mathbb{PD}$ give all necessary information to complete the task, but does not sound like PDDL, and does not describe all predicates needed to perform the task.
Natural $\mathbb{DD}$:
The Blocksworld game involves a set of blocks of different colors, which can be stacked on top of each other or placed on the table. The objective is to move the blocks from an initial configuration to a goal configuration using a series of legal moves. Legal moves in Blocksworld include: picking up a block from the table or from the top of another block, stacking a block onto the table, or stacking a block onto another block.
Natural $\mathbb{PD}$:
In this particular game, there are 4 blocks: a red block, a blue block, a green block, and a yellow block. At the start, the red block is on the table, the blue block is on top of the red block, the green block is on the table, and the yellow block is on top of the green block. The goal is to have the red block on the table, the green block on top of the red block, the yellow block on top of the green block, and the blue block on top of the yellow block.
A.3 MysteryBlocksWorld-100 PDDL
This section displays an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for MysteryBlocksWorld-100.
$\mathbb{DF}$:
(define (domain mystery_blocksworld) (:predicates (province ?x) (planet ?x) (harmony) (pain ?x) (craves ?x ?y)) (:action attack :parameters (?ob) :precondition (and (province ?ob) (planet ?ob) (harmony)) :effect (and (pain ?ob) (not (province ?ob)) (not (planet ?ob)) (not (harmony)))) (:action succumb :parameters (?ob) :precondition (pain ?ob) :effect (and (province ?ob) (harmony) (planet ?ob) (not (pain ?ob)))) (:action overcome :parameters (?ob ?underob) :precondition (and (province ?underob) (pain ?ob)) :effect (and (harmony) (province ?ob) (craves ?ob ?underob) (not (province ?underob)) (not (pain ?ob)))) (:action feast :parameters (?ob ?underob) :precondition (and (craves ?ob ?underob) (province ?ob) (harmony)) :effect (and (pain ?ob) (province ?underob) (not (craves ?ob ?underob)) (not (province ?ob)) (not (harmony)))))
$\mathbb{PF}$:
(define (problem mystery_blocksworld-p01) (:domain mystery_blocksworld) (:objects a b c d ) (:init (craves a b) (craves b c) (harmony) (planet c) (planet d) (province a) (province d) ) (:goal (and (craves a d) (craves c a) )) )
A.4 MysteryBlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
The following are example $\mathbb{DD}$ and $\mathbb{PD}$ of the Heavily Templated MysteryBlocksWorld-100. Text written in green demonstrates natural sounding text while text written in red demonstrates text that sounds the most like PDDL.
$\mathbb{DD}$:
I am playing with a set of objects. Here are the actions I can do Attack object Feast object from another object Succumb object Overcome object from another object I have the following restrictions on my actions: To perform Attack action, the following facts need to be true: Province object, Planet object, Harmony. Once Attack action is performed the following facts will be true: Pain object. Once Attack action is performed the following facts will be false: Province object, Planet object, Harmony. To perform Succumb action, the following facts need to be true: Pain object. Once Succumb action is performed the following facts will be true: Province object, Planet object, Harmony. Once Succumb action is performed the following facts will be false: Pain object. To perform Overcome action, the following needs to be true: Province other object, Pain object. Once Overcome action is performed the following will be true: Harmony, Province object, Object Craves other object. Once Overcome action is performed the following will be false: Province other object, Pain object. To perform Feast action, the following needs to be true: Object Craves other object, Province object, Harmony. Once Feast action is performed the following will be true: Pain object, Province other object. Once Feast action is performed the following will be false:, Object Craves other object, Province object, Harmony.
$\mathbb{PD}$:
As initial conditions I have that, object a craves object b, object b craves object c, harmony, planet object c, planet object d, province object a and province object d. My goal is to have that object a craves object d and object c craves object a.
Appendix B Prompts
For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-111:
Here is a game involving a table with blocks on it. {domain_description} {problem_description} Write the plan that would solve this problem. These are the available actions: (PICK-UP block): pick up a block from the table (PUT-DOWN block): put down a block on the table (STACK block1 block2): stack block1 onto block2 (UNSTACK block1 block2): unstack block1 from block2 Here is what the output should look like: (PICK-UP A) (STACK A B) (UNSTACK A B) (PUT-DOWN A)
For MysteryBlocksWorld-100, we use the following prompt:
Here is a game involving a table with blocks on it. {domain_description} {problem_description} Write the plan that would solve this problem. These are the available actions: (ATTACK object): attack object (SUCCUMB object): succumb (OVERCOME object1 object2): overcome object1 from object2 (FEAST object1 object2): feast object1 from object2 Here is what the output should look like: (ATTACK A) (OVERCOME A B) (FEAST A B) (SUCCUMB A)
Whenever possible, we asked the model to return the output in a JSON object for easier parsing.
Appendix C Results for Templated BlocksWorld-111
Metrics
Moderately Templated
Models Solvability Correctness
gemma-2-9b-it 0/111 0/111
gemma-2-9$\dagger$ - 5/111
gemma-2-27b-it 21/111 12/111
gemma-2-27$\dagger$ - 3/111
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gpt-3.5-turbo 14/111 4/111
gpt-4o-mini 9/111 5/111
gpt-4$\dagger$ - 7/111
gpt-4o 87/111 76/111
gpt-4$\dagger$ - 37/111
Heavily Templated
Models Solvability Correctness
gemma-2-9b-it 71/111 10/111
gemma-2-9$\dagger$ - 5/111
gemma-2-27b-it 90/111 89/111
gemma-2-27$\dagger$ - 9/111
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gpt-3.5-turbo 41/111 33/111
gpt-4o-mini 71/111 64/111
gpt-4$\dagger$ - 1/111
gpt-4o 99/111 99/111
gpt-4$\dagger$ - 32/111
Table 5: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on Moderately Templated and Heavily Templated BlocksWorld-111.
Table 5 displays results for Heavily Templated and Moderately Templated BlocksWorld-111.
Appendix D Results for Heavily Templated MysteryBlocksWorld-100
Table 6 displays the results for Heavily Templated MysteryBlocksWorld-100.
Models Solvability Correctness
Llama-3.1-8B 0/100 0/100
[…]
Llama-3.1-70B 0/100 0/100
[…]
gemma-2-9b-it 100/100 99/100
gemma-2-9$\dagger$ - 9/100
gemma-2-27b-it 99/100 98/100
gemma-2-27$\dagger$ - 0/100
gpt-3.5-turbo 4/100 0/100
gpt-4o-mini 36/100 5/100
gpt-4$\dagger$ - 0/100
gpt-4o 74/100 70/100
gpt-4$\dagger$ - 0/100
Table 6: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Heavily Templated MysteryBlocksWorld-100.
Appendix E Sample Model Output
The following is an example $\mathbb{DF}$ and $\mathbb{PF}$ that Llama-3.1-8B-Instruct gave. We can see that there are syntax errors, as well as semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$.
We give the model the following input $\mathbb{DD}$ and $\mathbb{PD}$:
Blocksworld is a puzzle domain where the aim is to move a set of blocks to achieve a specific goal configuration. The domain consists of a flat surface, often called the ’table’, and an arbitrary number of blocks. Every block can be stacked on top of another block or placed on the table. Blocks can be moved one at a time, and only the block at the top of a stack can be moved. The main actions available in Blocksworld are: 1. **Pick Up (Pick-Up B)**: Picks up block B from the table. 2. **Put Down (Put-Down B)**: Places block B onto the table. 3. **Stack (Stack B1 B2)**: Stacks block B1 on top of block B2. 4. **Unstack (Unstack B1 B2)**: Removes block B1 from the top of block B2. The goal is to transform an initial configuration of blocks into a desired target configuration using these actions. In this Blocksworld problem, we will consider a simple scenario with 2 blocks: Block1 and Block2. The initial state of the blocks is as follows: **Initial State**: - Block1 is on the table. - Block2 is on Block1. The goal is to achieve the following configuration: **Goal State**: - Block2 is on the table. - Block1 is on the table. To solve this problem, you will need to use the available actions to rearrange the blocks from the initial state to the goal state.
Which returns the following
Output $\mathbb{DF}$:
(define (domain blocksworld) (:requirements :strips :typing) (:types block) (:predicates (on-table?b - block) (on?b1 - block?b2 - block) (clear?b - block) ) (:action pick-up :parameters (?b - block) :preconditions (and (on-table?b) (clear?b)) :effects (and (not (on-table?b)) (not (clear?b)) (on-table?b)) ) (:action put-down :parameters (?b - block) :preconditions (and (not (on-table?b)) (clear?b)) :effects (and (on-table?b) (clear?b)) ) (:action stack :parameters (?b1 - block?b2 - block) :preconditions (and (on-table?b1) (clear?b2)) :effects (and (not (on-table?b1)) (not (clear?b2)) (on?b1?b2)) ) (:action unstack :parameters (?b1 - block?b2 - block) :preconditions (and (on?b1?b2) (clear?b1)) :effects (and (on-table?b1) (not (on?b1?b2)) (clear?b2)) ) )
Output $\mathbb{PF}$:
(define (problem blocksworld-problem) (:domain blocksworld) (:objects block1 block2 - block) (:init (on-table block1) (on block2 block1) (clear block1) (clear block2) ) (:goal (and (on-table block2) (on-table block1) (clear block1) (clear block2) ) ) )
Text in red displays errors outputted from the model. We can see that in the $\mathbb{DF}$ there are syntax errors (incorrect keyword “preconditions”) as well as semantic errors (incorrect predicates in preconditions and effects). For the output $\mathbb{PF}$ there are semantic errors, such as incorrect preconditions (a block cannot be clear and have another block on top of it) in the init section.
Environment LLM predicts? Natural Descriptions?
Zuo et al. (2024) fully-observed $\mathbb{PF}$ N
Zhang et al. (2024a) partially-observed $\mathbb{PF}$ N
Liu et al. (2023a) fully-observed $\mathbb{PF}$ N
Xie et al. (2023) fully-observed & partially-observed $\mathbb{PF}$ goal N
Lyu et al. (2023) fully-observed $\mathbb{PF}$ goal N
Zhang et al. (2024c) procedural texts $\mathbb{DF}$ action semantics N
Wong et al. (2023) partially-observed $\mathbb{DF}$ N
Guan et al. (2023) fully-observed $\mathbb{DF}$ $\mathbb{PF}$ goal N
Zhu et al. (2024) fully-observed $\mathbb{DF}$ action semantics N
Tang et al. (2024) partially-observed Python N/A
Silver et al. (2024) fully-observed plan N
Valmeekam et al. (2024) fully-observed plan N
Stechly et al. (2024) fully-observed plan N
This work fully-observed $\mathbb{DF}$ $\mathbb{PF}$ Y
Table 7: Comparison with related work.
Appendix F Related Works Comparison
Table 7 compares works related to this paper. We can see that other works as the LLM to predict either the plan, parts of PDDL files and other languages. We can also see that other works have mostly templated natural language descriptions, while this work uses both templated and natural descriptions.